Diffusion Morning Report03/25/11

Ted Berg tberg at stanford.edu
Fri Mar 25 06:43:45 PDT 2011


Good Morning All,
     Looking good this morning. Epi is back together and conditioning. 
LTO is going back together. Thermco poly1 had low/no dep-investigating. 
TEOS pump out for repair. Tylan 1 controller is coming back from repair. 
Thermco BPSG needs a test after repair. Ebeam compressor is dead. 
Thermco 2 no TLC flow. No other issues at this time.

					
	Date: 	3/25/11 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*down* 	*mfs out for repair* 	** 	** 	**
Tylan 2 	*up* 	** 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*down* 	*pump bad-out for repair* 	*.25/30* 	** 	**
Tylan 8-SiGe 	*up* 	** 	*.88/30* 	** 	**
Tylan 9-Poly 	*up* 	** 	*20.67/30* 	** 	**
Tylan 10-Nitride 	*up* 	** 	*9.3/30* 	** 	**
Tylan 11-LTO 	*down* 	*pull and clean* 	*0/30* 	** 	*1*
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*yellow* 	*no tlc flow* 	** 	** 	**
Thermco 3 	*future* 	** 	** 	** 	**
Thermco4 	*future* 	** 	** 	** 	**
Thermco Poly1 	*down* 	*NO DEP* 	** 	** 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*up* 	** 	** 	** 	**
Thermco bpsg 	*yellow* 	*needs test after repair* 	** 	** 	**
AMAT Epi 	*down* 	*conditioning* 	** 	** 	*2*
TEL Spa 	*up* 	** 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
Ebeam 	*down* 	*cryo compresssor dead* 	** 	** 	**
ZYGO 	*up* 	*
* 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	** 	*A on line 600 PSI* 	*changeover working 
well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20110325/e15f0374/attachment.html>


More information about the diffusion mailing list